Find the set of values of x for which
A . 3(2X+1)>5-2x
B.2x*2-7X+3>0
C..BOTH 3(2X+1)>5-2X AND 2X*2-7X+3>0

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

a.
3 * (2x + 1) > 5 - 2x
simplify to get:
6x + 3 > 5 - 2x
add 2x to both sides and subtract 3 from both sides to get:
8x > 2/8 which becomes x > .25.

b.
2x^2 - 7x + 3 > 0
set it equal to 0 to get:
2x^2 - 7x + 3 = 0
factor to get:
(2x-1) * (x-3) = 0
zero points are x = .5 and x = 3
2x^2 - 7x + 3 is less than zero between x = .5 and x = 3.
it is greater than zero when x is less than .5 and x is greater than 3.


to satisfy both inequalities, you hneed to have both x > .25 and x < .5 and x > 3.

your solution should be .25 < x < .5 or x > 3.

that can be seen in the following graph using desmos.com/calculator.



the area on the graph that is not shaded is your region of feasibility.

let me know if you have any questions.

theo


Answer by ikleyn(52832)

.

In your post, the input equations are written INCORRECTLY.

It is the same as, when driving a car, you miss the break pedal with the gas.